Four years ago, tom's age was 3/4 of the age he will be in 6 years. How old is he now?
Answer by macston(5194)

T=Tom's age
T-4=3/4(T+6)
T-4=3/4T+18/4 Add 4 to each side.
T=3/4T+34/4 Subtract 3/4T from each side.
1/4T=34/4 Multiply each side by 4.
T=34 ANSWER: Tom is currently 34 years old.
CHECK:
Four years ago Tom was 34-4=30
In six years he will be 34+6=40
3/4(40)=30
30=30